[Date Prev][Date Next][Thread Prev][Thread Next][Date Index][Thread Index]
Mars Pathfinder Update - September 19, 1997
- To: firstname.lastname@example.org
- Subject: Mars Pathfinder Update - September 19, 1997
- From: Ron Baalke <BAALKE@kelvin.jpl.nasa.gov>
- Date: Sun, 21 Sep 1997 0:18:48 GMT
- Old-X-Envelope-To: <email@example.com>
- Reply-To: firstname.lastname@example.org
- Resent-Date: Sat, 20 Sep 1997 20:19:31 -0400 (EDT)
- Resent-From: email@example.com
- Resent-Message-ID: <"5SoFzD.A.05G.sfGJ0"@mu.pair.com>
- Resent-Sender: firstname.lastname@example.org
MARS PATHFINDER MISSION STATUS
19 September 1997, 9:30 p.m. PDT
All Sol 76 uplink and downlink activities have been completed.
Following yesterday'ss downlink rain out, today's forecast once
again called for more soggy weather over the Deep Space Network
station in Canberra Australia. However, the rain held off today and
we were able to maintain a high data rate thought the entire Sol 76
downlink . Telemetry returned from Mars today indicated that our
Martian dwellers continue to be in good health after 2.5 months on
the surface on the red planet. The rover had another successful day
and images lost during yesterday's rain were retaken today and
received on the ground. The rover obtained images behind the rock
"Book Case" which is an area the lander is unable to see.
After acquiring the images the rover moved to the next target which
is the rock "Chimp". The rover stopped about a half meter to the
left of this rock because telemetry indicated a stall in the left
rear wheel which was possibly caused by a pebble being stuck in the
wheel. This has occurred in the past and the rover should be able to
clear this condition on its own.
The plan for tomorrow calls for completing the move to the rock
Chimp in preparation for placing the APXS on the rock.
More high rate data should be receive around 7:00 PM PDT tomorrow.
For further information on the Mars Pathfinder Mission, please call
our Mission Status Report line at 1-800-391-6654.